OpenStudy (anonymous):

Find f(x) and g(x) such that h(x)=(f*g)(x) h(x) = (7-6x)^2 Suppose that g(x)= 7-6x f(x)= ??

OpenStudy (anonymous):

f(x)=h(x)/g(x)= [(7-6x)^2]/(7-6x)=7-6x